OpenStudy (anonymous):

CORRECT ME 4 MEDAL [WORK IS SHOWN] Lacey and Rick have signed a contract to purchase a home. The closing date is June 27, and the buyer owns the property on the day of closing. The selling price of the home is $812,500. Lacey and Rick obtained a fixed-rate mortgage from a bank for $675,000 at 7.45% interest. The seller has already paid $14,878.15 in property taxes for the coming year. How much will Lacey and Rick owe in prorated expenses? $8,213.96 $65,165.65 $7,662.88 $25,900.76

OpenStudy (anonymous):

Jan 31 Feb28 March31 April 30 May 31 June 26 + --------- =177 days 365-177=188 days** 14878.15/365=40.76 40.76*188 = 7662.88 675000*0.00745/365=137.77 137.77*4=551.08 7662.88+551.08=8213.96

OpenStudy (anonymous):

My answer is right ! :D nevermind :3